OpenStudy (virnios):

Brad wants to buy flowers for his friend with $33. The daises are $1 each and the roses are $2 each. He buys 3 more daises than roses. How much did the roses cost? (I just need the equation and variable to solve for)

OpenStudy (anonymous):

each rose cost $2, so the cost of all roses Brad bought is equal $2*m, where m is number of roses. The same thing with daises.

OpenStudy (anonymous):

Brad spend $33 on all flowers, so we can write $33=$2*m + $1*n, where m is number of roses and n is number of daises.

OpenStudy (anonymous):

we also know that the number of roses is 3 less than number of daises or m=n - 3

OpenStudy (anonymous):

since we need to find the number of roses bought, the equation n=m+3 is better in this problem.

OpenStudy (anonymous):

now substitute m+3 into n in equation $33=$1*n + $2*m You get $33=$1(m+3) + $2*m Simplify. $33=$1m+$3+$2m. Combine like terms and solve further $33=$3+$3m

OpenStudy (virnios):

So it would be 10?

OpenStudy (anonymous):

yes

OpenStudy (virnios):

THANK YOU :)

OpenStudy (anonymous):

nope :) :)))